OpenStudy (anonymous):

Vendor sells candy to 500 customers at $15.00 per box. For each $.45 increase in the price of the candy, he would lose two customers. I need to determine what price he should see each box of candy to bring in the greatest total gross income.

OpenStudy (anonymous):

this is how much I have so far, X=boxes x(P)= X-2(P-15)/.45 P= price x(P)= 500-2(P-15)/.45 = 500-4.44(P-15) = 500-4.44P+66.6 = 566.6-4.44P